# Mantis: A Simulation-Grounded Foundation Model for Disease Forecasting

### v2

@@ Abstract
Infectious disease forecasting in novel outbreaks or low resource settings has been limited by the need for disease-specific data, bespoke training, and expert tuning. We introduce Mantis, a foundation model trained entirely on mechanistic simulations, which enables out-of-the-box forecasting across diseases, regions, and outcomes, even in settings with limited historical data. Mantis is built on over 400 million simulated days of outbreak dynamics spanning diverse pathogens, transmission modes, interventions, and surveillance artifacts. Despite requiring no real-world data during training, Mantis outperformed 39 expert-tuned models we tested across six diseases, including all models in the CDC’s COVID-19 Forecast Hub. Mantis generalized to novel epidemiological regimes, including diseases with hexld-out transmission mechanisms, demonstrating that it captures fundamental contagion dynamics. Critically, Mantis is mechanistically interpretable, enabling public health decision-makers to identify the latent drivers behind its predictions. Finally, Mantis delivers accurate forecasts at 8-week horizons, more than doubling the actionable range of most models, enabling proactive public health planning. Together, these capabilities position Mantis as a foundation for next-generation disease forecasting systems: general, interpretable, and deployable where traditional models fail.
@@ Introduction
Accurate forecasting of infectious disease dynamics is essential for effective public health response. Forecasts guide decisions around resource allocation, hospital staffing, vaccination campaigns, and non-pharmaceutical interventions such as school closures and social distancing [1, 2, 3]. Yet despite their importance, disease forecasts often fail in practice, particularly when they are most needed. During the first two years of the COVID-19 pandemic, the performance of many models in the CDC Forecast Hub was worse than a naïve baseline that simply projected the most recent value forward [4]. This highlights a key limitation of many operational systems: they struggle to provide accurate, actionable forecasts precisely when uncertainty is highest and decisions are most urgent.
@@
Several structural limitations contribute to this breakdown. First, most models require large volumes of disease-specific historical data, which are rarely available early in outbreaks. Second, they rely on extensive manual tuning by expert teams, limiting scalability and access in under-resourced health departments, such as those in rural counties or low- and middle-income countries. Third, they generalize poorly across diseases, locations, or targets, necessitating bespoke models for every context. In addition, their forecasting horizons are typically short: most models only predict 1–4 weeks ahead [5, 6, 4], far too narrow for real-world public health planning. Key decisions, such as ramping up hospital capacity, acquiring ventilators, or coordinating vaccination campaigns, often require 6–8 weeks of lead time. Without reliable long-range forecasts, these decisions are either delayed or made reactively, often at considerable cost [7].
Many high-performing models also lack interpretability. Neural networks can capture complex outbreak dynamics, but their predictions are often opaque [5], making them difficult to trust or act upon in high-stakes settings. This reflects a broader divide in the modeling landscape: mechanistic models offer transparency and domain alignment, but struggle with accuracy and adaptability. Machine learning models can deliver flexible, high-performance forecasts, but operate as black boxes. Public health decision-makers are not machine learning experts; they require forecasts that are not only accurate, but explainable in epidemiological terms. Without this, even good predictions may be dismissed or misunderstood. Bridging this gap requires tools that connect model behavior to domain-relevant mechanisms, such as transmission rates, reporting delays, or immunity waning. Enabling this kind of mechanistic interpretability could transform disease forecasting from a black-box tool into a trusted decision aid.
@@
Simulation-Grounded Neural Networks (SGNNs) provide a framework for overcoming these challenges [8]. SGNNs learn entirely from large-scale, highly realistic mechanistic simulations, using data generated by domain-specific simulators to train flexible neural networks. In epidemiology, this means generating synthetic outbreaks from mechanistic disease models that capture processes such as transmission, progression, immunity, seasonality, and intervention effects, then training models on these simulated trajectories so they can forecast real-world outbreaks they have never encountered before. By grounding learning in simulated systems that encode known mechanisms—whether physical, biological, or social—SGNNs enable models to generalize to real-world settings where data are limited, noisy, or missing. Crucially, they retain interpretability by linking predictions back to the underlying mechanistic simulations that informed them [8].
@@
We evaluate Mantis against 39 models, across forecasting tasks spanning six real-world diseases: seasonal influenza, early COVID-19, dengue, hepatitis B, smallpox, and scarlet fever. These evaluations cover multiple transmission modes, outbreak phases, surveillance systems, and historical eras. In every comparison we tested, including those against published, expert-tuned models and CDC Forecast Hub models, Mantis outperformed all baselines. At 8-week horizons, Mantis achieved lower relative error than the CDC’s ensemble of COVID-19 forecasts at 4 weeks, despite using no COVID-specific data or covariates in training.
Together, these capabilities establish Mantis as a general-purpose, interpretable forecasting system ready for real-world use and offering a scalable new blueprint for modeling in data-scarce or rapidly evolving settings.
To support deployment, we release Mantis as an open-source repository. Users can generate forecasts from a single historical time series without any machine learning expertise. 222github.com/carsondudley1/Mantis: Pretrained models, example data, and an interactive forecasting demo are available.
@@ Results
We assessed Mantis’s forecasting accuracy across a diverse benchmark suite of real-world tasks spanning six diseases: seasonal influenza, early COVID-19 [10], dengue [11], hepatitis B [12], smallpox [13], and scarlet fever [14]. For each task, we benchmarked Mantis against expert-tuned statistical and machine learning models, including published baselines and submissions to the CDC COVID-19 Forecast Hub [4] (see Methods).
Despite being trained entirely on synthetic data, Mantis consistently delivered state-of-the-art performance across all benchmarks we tested. Beyond raw accuracy, it enables reliable long-range forecasts up to 8 weeks ahead, generalizes out-of-the-box to novel diseases and epidemiological regimes, and provides mechanistic interpretability through back-to-simulation attribution.
Mantis Achieves State-of-the-Art Accuracy Across Diseases and Tasks
@@
For each disease, we evaluated performance across multiple spatial and temporal resolutions, including U.S. states and finer sub-state geographies, and at both the standard 4-week horizon [4] and a longer 8-week horizon. We compare Mantis against a range of published and re-implemented models, including top submissions to the CDC COVID-19 Forecast Hub, statistical baselines (naïve persistence and exponential smoothing), Chronos, a state-of-the-art time-series foundation model [15], and expert-tuned forecasting models previously developed for influenza.
Mantis outperformed every baseline across all diseases and tasks we tested (Figure 2a). This includes beating the CDC’s COVID-19 Forecast Hub ensemble, which combines dozens of expert-tuned models, even during volatile early-pandemic dynamics. We report forecasting improvement as the percent reduction in mean absolute error relative to the competing model for each task.
@@
Example forecast trajectories illustrate Mantis’s ability to match the timing and magnitude of real-world outbreaks with high fidelity (Figure 3). Forecasts include uncertainty intervals derived from the model’s predicted quantiles, enabling direct use in probabilistic planning. Additional forecasts of both real disease data and simulated Mantis training data can be found in the supplement.
Because Mantis is trained entirely on mechanistic simulations rather than historical observations, it requires no model fitting or fine-tuning for new targets. This enables true out-of-the-box deployment—an advantage over historically trained models that require task-specific data and manual calibration before producing reliable forecasts.
Figure 2: Mantis Delivers Strong and Consistent Forecasting Performance. (a) Mantis outperforms all baseline models across 39 forecasting tasks, including the CDC COVID-19 Forecast Hub ensemble. Bars represent percent improvement in absolute error relative to the best competing model, stratified by disease and task. The abbreviation SANP means same archticture as Mantis, no pretraining. ETS models are exponential smoothing models. (b) Including contextual covariates (e.g., using cases to predict hospitalizations) consistently improves Mantis’s accuracy across all forecast horizons, highlighting the model’s ability to integrate heterogeneous inputs to enhance generalization.
Covariate-Aware Mantis Improves Forecast Accuracy Across Horizons
@@
We evaluated two variants of Mantis to assess the impact of auxiliary input information on forecasting performance: a covariate-free version, which receives only the time series to be forecast (e.g., deaths), and a covariate-aware version, which also receives related signals such as cases and hospitalizations. This design mirrors real-world forecasting settings, where models may have access to upstream indicators with predictive value. (See Methods for training details for each model.)
@@
Across forecast horizons (2, 4, 6, and 8 weeks), the covariate-aware version consistently outperformed its covariate-free counterpart (Figure 2b), achieving lower mean absolute error across diseases and targets. These improvements were particularly pronounced at longer horizons, where upstream indicators (e.g., cases leading to hospitalizations, then deaths) provide early signals that improve downstream predictions.
@@
Mantis Enables Accurate Forecasts at Long Horizons
@@
Visual examples of 8-week forecasts show that Mantis captures outbreak peaks, rebounds, and multi-wave dynamics with high fidelity (Figure 3). Prediction intervals widen appropriately with time, maintaining calibrated uncertainty that reflects true forecast difficulty (for calibration results, see appendix D.1).
@@
We next demonstrate Mantis’s out-of-the-box forecasting capacity across both held-out transmission mechanisms and new input data modalities.
@@
First, we evaluated Mantis on diseases governed by transmission and progression mechanisms never represented in training simulations. For example, hepatitis B exhibits bloodborne transmission and chronic infection dynamics that are markedly different from the acute, self-limiting outbreaks used in pretraining. The Mantis training set included no models without recovery—all simulated diseases eventually transition out of the infectious state—and the longest average infectious period is just 10 days. In contrast, hepatitis B infection can persist for years. Despite this severe mismatch, Mantis outperformed statistical baselines trained directly on historical hepatitis B data (Fig. 3b, bottom right). These results demonstrate that Mantis can generalize across fundamental mechanistic boundaries—not just to new pathogens, but to qualitatively different epidemic regimes, such as chronic persistence versus outbreak spread.
@@
Next, we tested Mantis’s ability to generalize across data modalities. Although trained using cases, hospitalizations, and deaths, Mantis can incorporate new input types at inference time. We show that a Mantis model trained without ever seeing syndromic surveillance data (percent of outpatient visits that were for influenza-like illness) can still accurately forecast these syndromic indicators (Fig. 3a, right). This capability is especially valuable in settings where conventional case counts may be missing or delayed, but syndromic indicators are available in near real-time.
@@
In all settings we evaluated, Mantis outperforms comparison models trained directly on the task-specific data (Figure 2a). This level of generalization, across diseases, mechanisms, and modalities, underscores Mantis’s capacity as a true foundation model, and the advantages of scale afforded by simulation data. Rather than relying on domain-specific tuning, Mantis captures structural features of contagion that transfer across contexts, enabling immediate deployment in settings where other models would require retraining, expert tuning, or disease-specific data collection. Its ability to internalize general principles of transmission, rather than memorize specific patterns, makes it uniquely suited for forecasting in low-data and low-resource settings.
Mantis Provides Mechanistic Interpretability via Back-to-Simulation Attribution
Forecast accuracy alone is not sufficient in high-stakes public health settings. Decision-makers must be able to understand and act on model outputs, especially when choosing between interventions with different risk profiles. To support this, Mantis leverages back-to-simulation attribution (back2sim) [8], a form of mechanistic interpretability that identifies the underlying dynamics the model believes are driving a forecast.
This approach offers a different kind of explanation than post-hoc methods like SHAP or LIME [16, 17]. While those methods can assign importance scores to input features—telling you which data points were influential—they often fail to explain why in a way that is meaningful for scientific or policy decisions. Back-to-simulation attribution answers the central question: what mechanisms does model think are driving this outbreak?
The attribution process works by comparing a real-world outbreak to the library of scenarios Mantis was trained on. The procedure consists of the following steps:
  1. 1.
    First, an input time series (e.g., observed COVID-19 mortality) is processed by Mantis to create an internal summary of its key dynamic features.
  2. 2.
    This summary is then used to search the training dataset and identify the simulated outbreaks that are the most similar in character.
  3. 3.
    Critically, the model is trained to look beyond superficial resemblances in the data. The retrieved simulations are therefore not simply those with similar-looking curves, but rather those that Mantis has learned are driven by analogous underlying processes (e.g., similar transmission dynamics or immunity profiles).
  4. 4.
    Finally, the known, ground-truth parameters of this retrieved set of simulations are aggregated to construct a quantitative, mechanism-level explanation for the forecast.
This procedure reveals the model’s mechanistic belief about what is driving the observed data, offering a transparent, domain-aligned interpretation of its forecast.
We visualize the mechanistic beliefs inferred by Mantis using a held-out simulated trajectory for which the true latent parameters are known, allowing direct validation of back-to-simulation attribution (Fig. 4). In panel a, we show that the top-attributed simulations (gray lines) not only match the historical trajectory but also align closely with Mantis’s forecast, indicating that attribution selects plausible trajectories consistent with both past and projected dynamics. Panel b demonstrates that the parameters associated with these retrieved simulations track the ground-truth values across epidemic waves for key latent variables, including asymptomatic probability, case hospitalization rate, and hospitalization fatality rate. These results show that Mantis accurately recovers dynamic changes in underlying drivers. Panel c compares the attributed parameter distributions (blue) to the full simulation prior (gray), revealing sharp concentration around Mantis’s mechanistic beliefs and filtering out implausible settings. Finally, panel d confirms that attributed parameters cluster significantly closer to the true generative parameters than random samples from the prior, validating the precision and informativeness of Mantis’s internal representations. Together, these results demonstrate that Mantis not only forecasts accurately, but also provides interpretable, mechanism-level hypotheses about what is driving epidemic dynamics.
Figure 4: Mechanistic Interpretability via Back-to-Simulation Attribution. (a) Top-attributed simulations (gray) closely match the observed trajectory (red) and inform the forecast median and 90% confidence interval (blue), demonstrating temporal consistency and mechanistic plausibility. (b) Retrieved parameter estimates (black lines with 90% CI) for key latent variables—including asymptomatic transmission probability, hospitalization fatality rate, and case hospitalization rate—track true parameter values (red triangles) across epidemic waves, confirming the accuracy of back-to-simulation attribution. (c) Kernel density estimates of retrieved parameter distributions (blue) show strong sharpening relative to the simulation prior (gray), indicating Mantis’s ability to concentrate mechanistic belief around plausible drivers of observed dynamics. (d) Retrieved parameters cluster significantly closer to ground truth than randomly selected simulations from the prior, confirming the informativeness and specificity of attribution-based mechanistic explanations.

### v3

@@ Abstract
Infectious disease forecasting in novel outbreaks or low-resource settings is hampered by the need for disease-specific data, bespoke training, and expert tuning. We introduce Mantis, a foundation model trained entirely on mechanistic simulations, which enables out-of-the-box forecasting across diseases, regions, and outcomes, even in settings with limited historical data. We evaluated Mantis against 48 forecasting models across six diseases with diverse transmission modes, assessing both point forecast accuracy (mean absolute error) and probabilistic performance (weighted interval score and coverage). Despite using no real-world data during training, Mantis achieved lower mean absolute error than all models in the CDC’s COVID-19 Forecast Hub when backtested on early pandemic forecasts. Across all other diseases tested, including respiratory, vector-borne, and waterborne pathogens, Mantis consistently ranked in the top two models across all evaluation metrics. Notably, Mantis generalized to diseases with transmission mechanisms not represented in its training data, demonstrating that it captures fundamental contagion dynamics rather than memorizing disease-specific patterns. These capabilities position Mantis as a practical foundation for disease forecasting: general-purpose, accurate, and deployable where traditional models fail.
@@ Introduction
Accurate forecasting of infectious disease dynamics is fundamental to effective public health response. During the COVID-19 pandemic, forecasts touched nearly every aspect of daily life—from how scarce hospital beds were allocated to if restaurants could open and whether families could send their children to school [1]. Yet despite their importance, disease forecasts often fail in practice, particularly when they are most needed. During the first two years of the COVID-19 pandemic, the performance of many models in the CDC Forecast Hub was worse than a naïve baseline that simply projected the most recent value forward [2]. This highlights a key limitation of many operational systems: they struggle to provide accurate, actionable forecasts precisely when uncertainty is highest and decisions are most urgent.
@@
Several structural limitations contribute to this breakdown. First, most models require large volumes of disease-specific historical data, which are rarely available early in outbreaks. Second, they rely on extensive manual tuning by expert teams, limiting scalability and access in under-resourced health departments, such as those in rural counties or low- and middle-income countries. Third, they generalize poorly across diseases, locations, or targets, necessitating new models for every context. In addition, their forecasting horizons are typically short: most models only predict 1–4 weeks ahead [2, 3, 4], far too short-term for real-world public health planning. Key decisions, such as ramping up hospital capacity, acquiring ventilators, or coordinating vaccination campaigns, often require 6–8 weeks of lead time. Without reliable long-range forecasts, these decisions are either delayed or made reactively, often at considerable cost [5].
Typically, projections beyond 4 weeks are restricted to scenario analysis, which explores hypothetical outcomes conditional on specific policy interventions or behavioral changes [6, 7]. However, unconditional forecasts at longer horizons remain valuable for public health planning. We define forecasts as unconditional projections of expected outcomes, versus scenario analyses that specify hypothetical interventions. By this definition, temporal horizon does not determine whether a projection is a forecast or a scenario—an 8-week unconditional projection is a forecast, while a 2-week conditional projection is still scenario analysis. Public health decisions such as hospital capacity expansion and vaccination campaigns often require 6-8 weeks of lead time, motivating the development of models capable of reliable longer-range forecasting without any policy assumptions.
@@
Addressing these challenges (limited historical data, poor cross-disease generalization, restricted forecast horizons, and the need for extensive expert tuning) requires a different approach to model training. Simulation-Grounded Neural Networks (SGNNs) provide such a framework [8]. SGNNs learn entirely from large-scale, highly realistic mechanistic simulations, using data generated by domain-specific simulators to train flexible neural networks. In epidemiology, this means generating synthetic outbreaks from mechanistic disease models that capture processes such as transmission, progression, immunity, seasonality, and intervention effects, then training models on these simulated trajectories so they can forecast real-world outbreaks they have never encountered before. By grounding learning in simulated systems that encode known mechanisms—whether physical, biological, or social—SGNNs enable models to generalize to real-world settings where data are limited, noisy, or missing.
@@
We evaluate Mantis against 48 models across forecasting tasks spanning six real-world diseases: seasonal influenza, early COVID-19, dengue, hepatitis B, smallpox, and scarlet fever. These evaluations cover multiple transmission modes, outbreak phases, surveillance systems, and historical eras. We assessed both point forecast accuracy and probabilistic performance. When benchmarked against CDC Forecast Hub models on early pandemic COVID-19 forecasts, Mantis achieved the lowest mean absolute error despite using no COVID-specific data or covariates in training and was tied for the second lowest weighted interval score. Across all other diseases—including those with transmission mechanisms absent from its training data—Mantis consistently ranked among the top two models on all evaluation metrics, demonstrating robust generalization to diverse epidemiological settings.
These capabilities establish Mantis as a general-purpose forecasting system offering a scalable solution for modeling in data-scarce or rapidly evolving settings. To support deployment, we release Mantis as an open-source repository. Users can generate forecasts from a single historical time series without any machine learning expertise. 222github.com/carsondudley1/Mantis: Pretrained models, example data, and an interactive forecasting demo are available. Mantis is currently contributing forecasts to a new local-scale forecasting hub within Michigan for the 2025-2026 respiratory virus season and will be submitted to the CDC Forecast Hub to support real-time public health decision-making.
@@ Results
We assessed Mantis’s forecasting accuracy across a set of real-world forecasting tasks spanning six diseases: seasonal influenza [10, 11], early COVID-19 [12], dengue [13], hepatitis B [14], smallpox [15], and scarlet fever [16]. For each task, we benchmarked Mantis against expert-tuned statistical and machine learning models, including published baselines and submissions to the CDC COVID-19 Forecast Hub [2] (see Methods). We evaluated both point forecast accuracy using mean absolute error (MAE) and probabilistic forecasting performance using weighted interval score (WIS) and coverage metrics (e.g., how often did the 95% CI actually contain the true value).
Mantis Achieves Competitive Performance Across Diseases and Tasks
For each disease, we evaluated performance across multiple spatial and temporal resolutions, including U.S. states and smaller regions within states. We evaluated both point forecast performance (mean absolute error, MAE) and probabilistic metrics (weighted interval score, WIS; 50% and 95% coverage). WIS is the standard probabilistic evaluation metric for disease forecasting [17], providing a comprehensive assessment that accounts for both forecast accuracy and uncertainty calibration across predicted quantiles. Following the evaluation framework established in [2], we report relative MAE and relative WIS, where metrics are normalized by the performance of a naïve baseline that forecasts the current value forward with uncertainty intervals derived from historical variance (see Methods). Relative metrics offer two key advantages over absolute error: they provide an interpretable measure of the information contributed by a model beyond simple persistence, and they enable meaningful comparisons across different forecast targets and geographic scales, where absolute case or death counts may vary by orders of magnitude.
@@
We compare Mantis against a range of published and re-implemented models, including top submissions to the CDC COVID-19 Forecast Hub, statistical baselines (exponential smoothing and SARIMA), LSTM neural networks retrained at each forecast point, and Chronos, a state-of-the-art time-series foundation model [18].
Mantis achieved competitive performance across diseases and metrics (Table 1). For early pandemic COVID-19 forecasts at the state level, Mantis achieved the lowest MAE among evaluated models, including the CDC Forecast Hub ensemble that combines dozens of expert-tuned submissions. The CDC ensemble achieved slightly better WIS (0.61 vs 0.63), while UMass-MechBayes achieved marginally better 95% coverage (0.93 vs 0.91), demonstrating that multiple approaches can excel on different aspects of probabilistic forecasting. Although results have shown that some general time series forecasting foundation models can confer reasonable results for infectious disease forecasting [19], we did not find that any we tested gave competitive performance with Mantis or our other baseline models, so we did not include them in our analysis.
Across the five additional diseases tested—influenza hospitalizations, dengue, smallpox, scarlet fever, and hepatitis B—Mantis consistently ranked first or second on both point forecast accuracy (MAE) and probabilistic performance (WIS).
@@
Example forecast trajectories illustrate Mantis’s ability to match the timing and magnitude of real-world outbreaks across diverse epidemiological settings (Figure 3). Forecasts include uncertainty intervals derived from the model’s predicted quantiles, enabling direct use in probabilistic planning. Because Mantis is trained entirely on broad mechanistic simulations rather than historical observations, it requires no model fitting or fine-tuning for new targets.
Table 1: Forecasting results across diseases and models. For each metric, the best model is highlighted in green and bold, and the second-best is blue and underlined.
Disease Model
Relative MAE
Relative WIS
50% coverage
95% coverage
State-level COVID Mantis 0.65 0.63 0.51 0.91
CDC ensemble 0.66 0.61 0.53 0.90
UMass-MechBayes 0.67 0.63 0.56 0.93
GT-DeepCOVID 0.82 0.75 0.41 0.84
UCSD_NEU-DeepGLEAM 0.78 0.83 0.70 0.91
SARIMA 1.08 1.10 0.44 0.81
LSTM 1.37 1.54 0.33 0.70
Local-scale flu hosp. Mantis 0.85 0.86 0.51 0.92
ETS 1.15 1.32 0.39 0.78
LSTM 1.21 1.33 0.40 0.79
SARIMA 1.10 1.00 0.47 0.78
Dengue cases Mantis 0.75 0.80 0.47 0.90
ETS 0.99 1.04 0.45 0.90
LSTM 1.34 1.62 0.37 0.73
SARIMA 0.94 1.58 0.73 0.92
Smallpox cases Mantis 0.83 0.84 0.45 0.87
ETS 1.00 1.06 0.44 0.86
LSTM 1.31 1.52 0.33 0.71
SARIMA 0.95 1.04 0.53 0.88
ILI syndromic Mantis 0.83 0.88 0.46 0.89
ETS 1.00 1.05 0.43 0.83
LSTM 1.19 1.30 0.41 0.80
SARIMA 1.03 0.97 0.45 0.88
Scarlet Fever cases Mantis 0.80 0.81 0.47 0.91
ETS 0.89 0.95 0.46 0.89
LSTM 0.87 1.04 0.36 0.80
SARIMA 1.05 1.08 0.56 0.90
Hepatitis B cases Mantis 0.85 0.84 0.47 0.90
ETS 0.86 0.89 0.46 0.88
LSTM 0.82 1.07 0.37 0.78
SARIMA 0.86 0.84 0.58 0.91
Covariates Improve Forecast Accuracy Across Horizons
@@
We evaluated two variants of Mantis to assess the impact of additional information on forecasting performance: a covariate-free version, which receives only the time series to be forecast (e.g., deaths), and a covariate-aware version, which also receives related signals (e.g., cases or hospitalizations). This design mirrors real-world forecasting settings, where models may have access to upstream indicators with predictive value. (See Methods for training details for each model.)
@@
Across forecast horizons (2, 4, 6, and 8 weeks), the covariate-aware version consistently outperformed its covariate-free counterpart (Figure 2a), achieving lower mean absolute error across diseases and targets. These improvements were particularly pronounced at longer horizons, where upstream indicators (e.g., cases leading to hospitalizations, then deaths) provide early signals that improve downstream predictions.
@@
Figure 2: Covariate integration improves accuracy. Mantis maintains calibrated uncertainty across forecast horizons. (a) Including covariates (e.g., using cases to predict hospitalizations) consistently improves Mantis’s accuracy across all forecast horizons. Relative MAE shown for COVID-19 mortality forecasts with (blue) and without (orange) hospitalization covariates across 2, 4, 6, and 8-week horizons. (b) Mantis’s prediction interval widths increase appropriately with forecast horizon, reflecting growing uncertainty in longer-range predictions. The 50% confidence interval width (blue) grows from approximately 50 to 80 deaths, while the 95% interval width (orange) expands from approximately 110 to 180 deaths over 1 to 4 weeks ahead when forecasting COVID-19 mortality. This systematic widening demonstrates the model becoming appropriately less confident as prediction horizon increases.
Mantis Enables Accurate Forecasts at Longer Horizons
@@
Visual examples of 8-week forecasts are shown in figure 3b. Prediction intervals widen with time (figure 2b), maintaining calibrated uncertainty that reflects forecast difficulty (calibration results in table 1).
@@
We next evaluated Mantis’s out-of-the-box forecasting capacity across both held-out transmission mechanisms and new input data modalities.
@@
First, we benchmarked Mantis on diseases governed by transmission and progression mechanisms never represented in training simulations. For example, hepatitis B exhibits bloodborne transmission and chronic infection dynamics that are markedly different from the acute, self-limiting outbreaks used in pretraining. The Mantis training set included no models without recovery—all simulated diseases eventually transition out of the infectious state—and the longest average infectious period is just 10 days. In contrast, hepatitis B infection can persist for years. Despite this fundamental mechanistic mismatch, Mantis achieved competitive performance on hepatitis B forecasts, ranking second in MAE (0.85 vs 0.82) and achieving the best WIS (0.84) among all models (Table 1). These results demonstrate that Mantis can generalize across fundamental mechanistic boundaries—not just to new pathogens, but to qualitatively different epidemic regimes, such as chronic persistence versus outbreak spread.
@@
Next, we tested Mantis’s ability to generalize across data modalities. Although trained using cases, hospitalizations, and deaths, Mantis can incorporate new input types at inference time. We evaluated Mantis on syndromic surveillance data (percent of outpatient visits for influenza-like illness), a data type never represented in training simulations. Despite this distribution shift, Mantis achieved the lowest MAE (0.83) and WIS (0.88) on ILI forecasting (Table 1), demonstrating robust generalization to novel surveillance indicators. This capability is especially valuable in settings where conventional case counts may be missing or delayed, but syndromic indicators are available in near real-time.
@@
Across all seven disease forecasting tasks, spanning respiratory, vector-borne, bloodborne, and syndromic targets, Mantis consistently ranked first or second on both point and probabilistic metrics (Table 1). This level of generalization, across diseases, mechanisms, and modalities, demonstrates Mantis’s capacity as a foundation model and the advantages of scaling afforded by simulation data. Rather than relying on domain-specific tuning, Mantis captures structural features of contagion that transfer across contexts, enabling immediate deployment in settings where other models would require retraining, expert tuning, or disease-specific data collection.